# Lexicon3D: Probing Visual Foundation Models for Complex 3D Scene Understanding

### v1

@@ B.3 Additional Qualitative Results
Figure B: Visualization of partial scene registration results. The family of generative models StableDiffusion and StableVideoDiffusion obtains superior performance. Also, video encoders like V-JEPA and StableVideoDiffusion have better geometric understanding capability than image encoders.
@@ Appendix C Limitations and Future Work
Model capacities not strictly identical or comparable.   Our evaluation focuses on seven vision foundation models due to their availability and common use in recent work. Consequently, all our experiments are based on publicly available checkpoints released by the project owners. Although we have attempted to choose models with similar capacities, achieving strictly identical backbone architectures was not possible without re-training all the baselines ourselves. However, such experiments require an enormous amount of computational resources that we cannot afford.
Our evaluation focuses on indoor scenarios.   Recent literature often separates the study of indoor scene perception and reasoning from outdoor scenarios, which are often relevant to autonomous driving or robotics applications. Outdoor scenarios present different challenges compared to indoor scenes. Lexicon3D focuses its evaluation solely on indoor scenes. While this is a valid choice considering that most scene-level multi-modal benchmarks are still based on indoor scenes, it is not comprehensive. Outdoor scenarios contain large ego-movement speeds and many more dynamic moving objects than indoor scenes. Evaluating these scenes will likely lead to unique observations, and we consider this a direct future direction.
We adopt the most straightforward approach to probing.   To evaluate the capabilities of the visual foundation models, we freeze their parameters and only tune the linear or shallow probing head. This approach allows us to analyze the pretrained methods’ capabilities without altering their models through the finetuning process. While we argue that probing the frozen encoder provides the most accurate understanding of these models, we acknowledge that the ability to quickly adapt to new tasks with finetuning is also an important aspect of an encoder. However, finetuning these large-scale models, which often have close to billion-level parameters, requires a significant amount of time and computational resources. We leave this study for future work.
@@ Appendix D Societal Impact
We anticipate a potential positive social impact from our work. Lexicon3D represents one of the first steps towards a comprehensive understanding of large-scale visual foundation models in real-world 3D scene analysis and reasoning. This understanding could lead to the development of more robust and efficient scene encoding systems, benefiting a wide range of applications, including autonomous driving, virtual reality, household robots, and multi-modal chatbots. Ultimately, this could contribute to a more inclusive, efficient, and safer world, where technology understands and adapts to the diverse ways humans perceive and navigate their environments.
Potential negative societal impact.   We do not see a direct negative societal impact on our work. Indirect potential negative impact involves misusing strong scene encoding foundation models for surveillance or virtual reality. We believe it is crucial for researchers to proactively consider these concerns and establish guidelines to ensure the responsible usage of these models.

### v2

@@ B.4 Additional Qualitative Results
Figure B: Visualization of partial scene registration results. The StableDiffusion and StableVideoDiffusion family of generative models receives superior performance. In addition, video encoders such as V-JEPA and StableVideoDiffusion have better geometric understanding capability than image encoders.
@@ Appendix C Limitations and Future Work
Model capacities are not strictly identical or comparable.   Our evaluation focuses on seven vision foundation models due to their availability and common use in recent work. Consequently, all our experiments are based on publicly available checkpoints. Although we have attempted to choose models with similar capacities, achieving strictly identical backbone architectures was not possible without re-training all the baselines ourselves. However, such experiments require an enormous amount of computational resources that we cannot afford.
Our evaluation focuses on indoor scenarios.   Recent literature often separates the study of perception and reasoning of indoor scenes from outdoor scenarios, which are often relevant to autonomous driving or robotics applications. Outdoor scenarios present different challenges compared to indoor scenes. Lexicon3D focuses its evaluation solely on indoor scenes. While this is a valid choice considering that most scene-level multimodal benchmarks are still based on indoor scenes, it is not comprehensive. Outdoor scenarios contain large ego-movement speeds and many more dynamic moving objects than indoor scenes. Evaluating these scenes will likely lead to unique observations, and we consider this a direct future direction.
We adopt the most straightforward approach to probing.   To evaluate the capabilities of the visual foundation models, we freeze their parameters and only tune the linear or shallow probing head. This approach allows us to analyze the capabilities of the pretrained methods without altering their models through the finetuning process. Although we argue that probing the frozen encoder provides the most accurate understanding of these models, we acknowledge that the ability to quickly adapt to new tasks with finetuning is also an important aspect of an encoder. However, finetuning these large-scale models, which often have close to billion-level parameters, requires a significant amount of time and computational resources. We leave this study for future work.
@@ Appendix D Societal Impact
We anticipate a potential positive societal impact from our work. Lexicon3D represents one of the first steps towards a comprehensive understanding of large-scale visual foundation models in real-world 3D scene analysis and reasoning. This understanding could lead to the development of more robust and efficient scene encoding systems, which benefit a wide range of applications, including autonomous driving, virtual reality, household robots, and multimodal chatbots. Ultimately, this could contribute to a more inclusive, efficient, and safer world, where technology understands and adapts to the diverse ways humans perceive and navigate their environments.
Potential negative societal impact.   We do not see a direct negative societal impact on our work. Indirect potential negative impact involves misusing strong scene-encoding foundation models for surveillance or virtual reality. We believe that it is crucial for researchers to proactively consider these concerns and establish guidelines to ensure responsible usage of these models.